OpenStudy (anonymous):

Algebra 2 Word Problem: Set up using 3 variables & 3 equations: "A friend e-mails you the results of a recent high school swim meet. The e-mail states that 24 individuals placed, earning a combined total of 53 points. First place earned 3 points, second place earned 2 points, and third place earned 1 point. There were as many first-place finishers as second and third-place finishers combined." A: Write a system of THREE equations that represents how many people finished in each place. B: How many swimmers finished in first place, second and third?

OpenStudy (wolf1728):

When you say 24 individuals placed do you mean 24 people finished first second or third?

OpenStudy (wolf1728):

f + s + t = 24 3*f + 2*s + t = 53 f = (s + t)

OpenStudy (wolf1728):

First Place 12 Second Place 5 Third Place 7 12*3 = 36 5*2 =10 plus 7 sums to 53 12+5+7 sums to 24 participants 12 (first place) = (5 second place and 7 third place) That seems to be it.